What was found
LEFT PACK OUTPUT TEMP IS STILL UNSTABLE, RIGHT PACK IS STABLE. LEFT VARIES +/- 6 DEGREES FROM RIGHT PACK REFERENCE TEMP. GROUNDED - YES. PERFORMED DIVERGENT PACK OUTLET TEMPERATURES NOTED BETWEEN LEFT AND RIGHT PACK FOUND T475 RT MIX MANIFOLD TEMP SENSOR HIGH IN REF TO FIM 21-53 TASK 819. NO HELP. LT PACK TEMP CONTROLLER REMOVED AND LT MIX MANIFOLD TEMP SENSOR D3868 DISCONNECTED. CONTINUITY CHECK OF WIRING GOOD. PACK TEMP CONTROLLER RE-INTEMP SENSOR T472 REMOVED AND REPLACED REF AMM 21-51-51. ZONE TEMP CONTROLLER BITE TEST GOOD REF FIM 21-62 TASK 801. OPS CHECK GOOD.